# SPEECH RECOGNITION ON TV SERIES WITH VIDEO-GUIDED POST-ASR CORRECTION

### v2

@@ Abstract
Index Terms—  post-ASR correction, video-audio, video-large multimodal model
@@ 1 Introduction
Automatic Speech Recognition (ASR) has achieved remarkable success with the advent of deep learning, driving advancements in applications such as conversational AI, media transcription, and assistive technologies [1]. Despite their success, ASR systems still struggle in complex real-world scenarios, particularly in noisy environments [2], domain-specific conversations [3], and multimedia content such as TV series or movies [4]. These scenarios often involve overlapping speech, domain-specific terminology, and contextual dependencies, making it difficult for ASR systems to rely solely on audio signals to achieve optimal performance.
To address the above challenges, leveraging multimodal contextual information has emerged as a promising direction [5]. However, the existing works focusing on utilizing image information to correct the ASR output [6] lack full exploitation of video information [7]. ASR on TV series [8, 9] is particularly important for accessibility and engagement. In particular, video serves as a powerful complementary modality, providing rich contextual cues that speech alone cannot convey. For example, the video-based context, such as scene elements, character movements, and long-range dependencies inherent in TV series and movies, can significantly aid in maintaining coherence and provide more useful information for recognizing named entities.
@@
We evaluate our approach on a multimodal TV series dataset Violin [10] with state-of-the-art self-supervised speech models (wav2vec 2.0 [11], HuBERT [12], WavLM [13]), and Conformer [14]. The experiments demonstrate that our method can significantly reduce the Word Error Rate (WER) by providing necessary visual context, yielding a relative 20.75% improvement on the Violin dataset [10] for WavLM.
@@ 2 Related Work
Recent post-ASR correction [15] aims at correcting the generative errors of a pretrained ASR model using LLMs, with training-free strategy [16] or light additional training [17]. However, real-world ASR tasks often occur in multimodal contexts, where additional information from video can enhance error correction. Audio-visual models [6, 7, 18] improve recognition by jointly encoding visual and audio features. In contrast, the role of video context in post-ASR correction remains underexplored, despite its potential to mitigate errors that persist after decoding. While these models have demonstrated strong performance, an important research gap remains: the potential role of visual information in further improving ASR performance. VLMMs [19, 20, 21] have excelled in tasks such as video understanding, captioning, and video-based question answering, but their application to ASR correction remains unexplored. Therefore, different from the existing work, we propose a novel training-free multimodal post-ASR correct framework, which integrates visual context from video data to improve the quality of ASR outputs.
Fig. 1: Overview of the proposed Video-Guided Post-ASR Correction (VPC) method. The framework integrates video-based contextual information to refine ASR outputs.
@@ 4.1 ASR Generation
Despite the compelling success recent ASR systems have achieved in several datasets [22, 23, 24], they still struggle with named entity misrecognition errors, phonetic approximation errors, and homophone errors. For example, when transcribing TV series dialogues, it may misrecognize Joey Tribbiani as Joey Tribbyany. This error occurs because the model lacks sufficient world knowledge to correctly infer the proper spelling of named entities, especially those that are uncommon or do not frequently appear in the training data. Therefore, we devised a video-guided post-ASR correction method to correct the output for ASR models. In this stage, we feed the ASR model with audio and get the corresponding generated transcript $\hat{Y}$.
@@ 4.2.1 Video-based Contextual Information Extraction
Given VLMMs’ impressive performance in video understanding, we resort to VLMM to capture the rich semantic context (e.g., scenes, actions, and characters) in video. Specifically, we extract the fine-grained information from the video with a Question-Answering (QA) format. We devised two questions to fully mine the information from the video: 1) TV Show Recognition and 2) Fine-grained Video Description. The potential reason we devised these two questions is: 1) the TV show name can be utilized to retrieve knowledge related to the TV show, such as character names, and the QA-based information extraction method could mine this kind of information from videos; 2) ASR models sometimes generate errors that are unrelated to the video content or topic, and fine-grained video description could be used to correct such errors.
@@ 5 Experiments
5.1 Experimental Setup
5.1.1 Dataset.
We evaluate our approach on the Violin dataset [10], which provides multimodal content including audio, video, and subtitles from TV series, movies, and YouTube clips. While the dataset provides diverse clips, some contain only background music without speech, while others include English subtitles but are spoken in different languages.
@@ 5.1 Experimental Setup
5.1.2 Implementation Details.
To assess our VPC’s effectiveness in correcting transcription, we selected wav2vec 2.0, HuBERT and WavLM, and Conformer as our ASR models. All models are pretrained on Librispeech-960h [24]. We finetuned wav2vec 2.0, HuBERT and WavLM with CTC loss, and finetuned Conformer with RNN-T loss on the 72-hour training set. In our VPC method, we apply VideoLLaMA2 [19] as our VLMM, and GPT-4o as our LLM. For the evaluation metric, we use WER.
@@ 5.2 Results Analysis
Fig. 2: Comparison between results generated by WavLM and its corresponding variant augmented with VPC.
@@ 5.3 Case Study
We further show a case study about the comparison between results generated by WavLM and its corresponding variant augmented with our method. In this sample, our proposed model revises the error words “a be hi hat” into the correct word “a beehive”. This is because the video content provides additional information to facilitate the LLM to revise the potential error by reasoning. This further verifies the effectiveness of our method.

### v3

# Yang Zhang Jing Hansen

@@ Abstract
keywords:
post-ASR correction, video-audio, video-large multimodal model
@@ 1 Introduction
Automatic Speech Recognition (ASR) has achieved remarkable success with the advent of deep learning, driving advancements in applications such as conversational AI, media transcription, and assistive technologies [DBLP:conf/icassp/GravesMH13]. Despite their success, ASR systems still struggle in complex real-world scenarios, particularly in noisy environments [DBLP:journals/ijst/DuaAD23], domain-specific conversations [DBLP:conf/um/CaoGCSPRKMPD23], and multimedia content such as TV series or movies [DBLP:conf/jauti/BecaASR18]. These scenarios often involve overlapping speech, domain-specific terminology, and contextual dependencies, making it difficult for ASR systems to rely solely on audio signals to achieve optimal performance.
To address the above challenges, leveraging multimodal contextual information has emerged as a promising direction [DBLP:conf/icassp/NiHC00L024]. However, the existing works focusing on utilizing image information to correct the ASR output [DBLP:conf/interspeech/GabeurSN0AS22] lack full exploitation of video information [DBLP:conf/cvpr/SeoNS23]. ASR on TV series [DBLP:conf/interspeech/GuinaudeauGS10, DBLP:conf/icassp/KorbarHZ24] is particularly important for accessibility and engagement. In particular, video serves as a powerful complementary modality, providing rich contextual cues that speech alone cannot convey. For example, the video-based context, such as scene elements, character movements, and long-range dependencies inherent in TV series and movies, can significantly aid in maintaining coherence and provide more useful information for context-dependent disambiguation.
@@
We evaluate our approach on a multimodal TV series dataset Violin [DBLP:conf/cvpr/LiuCCGYYL20] with state-of-the-art self-supervised speech models (wav2vec 2.0 [DBLP:conf/nips/BaevskiZMA20], HuBERT [DBLP:journals/taslp/HsuBTLSM21], WavLM [DBLP:journals/jstsp/ChenWCWLCLKYXWZ22]), and Conformer [DBLP:conf/interspeech/GulatiQCPZYHWZW20]. The experiments demonstrate that our method can significantly reduce the Word Error Rate (WER) by providing necessary visual context, yielding a relative 20.75% improvement on the Violin dataset [DBLP:conf/cvpr/LiuCCGYYL20] for WavLM.
@@
Figure 1: Overview of the proposed Video-Guided Post-ASR Correction (VPC) method. The framework integrates video-based contextual information to refine ASR outputs.
@@ 2 Related Work
Recent post-ASR correction [DBLP:journals/corr/abs-2310-13013] aims at correcting the generative errors of a pretrained ASR model using LLMs, with training-free strategy [DBLP:journals/corr/abs-2505-24347] or light additional training [DBLP:conf/emnlp/RadhakrishnanYK23]. However, real-world ASR tasks often occur in multimodal contexts, where additional information from video can enhance error correction.
While traditional Audio-Visual Speech Recognition (AVSR) methods, such as AV-HuBERT, have demonstrated success in improving ASR robustness, they primarily rely on low-level sensory fusion [DBLP:conf/interspeech/GabeurSN0AS22, DBLP:conf/cvpr/SeoNS23, DBLP:conf/iclr/ShiHLM22_avhubert]. These models focus on extracting lip-reading cues and facial movements, which require high-resolution, perfectly aligned face tracks to function effectively.
However, in the complex environment of TV series, such fine-grained visual data is often unavailable due to off-screen speakers, wide shots, or low lighting; at the same time the higher-level context information are overlooked.
In contrast, the role of video context in post-ASR correction remains underexplored, despite its potential to mitigate errors that persist after decoding. While these models have demonstrated strong performance, an important research gap remains: the potential role of visual information in further improving ASR performance. VLMMs [DBLP:journals/corr/abs-2406-07476, DBLP:conf/acl/0001RKK24, DBLP:conf/emnlp/LinYZCNJ024] have excelled in tasks such as video understanding, captioning, and video-based question answering, but their application to ASR correction remains unexplored. Therefore, different from the existing work, we propose a novel training-free multimodal post-ASR correct framework, which integrates visual context from video data to improve the quality of ASR outputs.
@@ 4.1 ASR Generation
Despite the compelling success recent ASR systems have achieved in several datasets [DBLP:conf/icassp/KahnRZKXMKLCFLS20, DBLP:journals/corr/abs-1207-0580, DBLP:conf/icassp/PanayotovCPK15], they still struggle with context-dependent disambiguation, phonetic approximation errors, and homophone errors. For example, when transcribing TV series dialogues, it may misrecognize Joey Tribbiani as Joey Tribbyany. This error occurs because the model lacks sufficient world knowledge to correctly infer the proper spelling of ambiguous words that is context-dependent, especially those that are uncommon or do not frequently appear in the training data. Therefore, we devised a video-guided post-ASR correction method to correct the output for ASR models. In this stage, we feed the ASR model with audio and get the corresponding generated transcript $\hat{Y}$.
@@ 4.2.1 Video-based Contextual Information Extraction
Given VLMMs' impressive performance in video understanding, we resort to VLMM to capture the rich semantic context (e.g., scenes, actions, and characters) in video. Specifically, we extract the fine-grained information from the video with a Question-Answering (QA) format. We devised two questions to fully mine the information from the video: 1) TV Show Recognition and 2) Fine-grained Video Description. The potential reason we devised these two questions is: 1) the TV show name can be utilized to retrieve knowledge related to the TV show, such as character names, and the QA-based information extraction method could mine this kind of information from videos; 2) ASR models sometimes generate errors that are unrelated to the video content or topic, and fine-grained video description could be used to correct such errors.
@@ 5 Experiments
5.1 Dataset.
We evaluate our approach on the Violin dataset [DBLP:conf/cvpr/LiuCCGYYL20], which provides multimodal content including audio, video, and subtitles from TV series, movies, and YouTube clips. While the dataset provides diverse clips, some contain only background music without speech, while others include English subtitles but are spoken in different languages.
@@
5.2 Implementation Details.
To assess our VPC's effectiveness in correcting transcription, we selected wav2vec 2.0, HuBERT and WavLM, and Conformer as our ASR models. All models are pretrained on Librispeech-960h [DBLP:conf/icassp/PanayotovCPK15]. We finetuned wav2vec 2.0, HuBERT and WavLM with CTC loss, and finetuned Conformer with RNN-T loss on the 72-hour training set. In our VPC method, we apply VideoLLaMA2 [DBLP:journals/corr/abs-2406-07476] as our VLMM, and GPT-4o as our LLM. For the evaluation metric, we use WER.
We considered including state-of-the-art joint Audio-Visual ASR (AV-ASR) models, such as AV-HuBERT [DBLP:conf/iclr/ShiHLM22_avhubert], as baselines. However, our preliminary experiments of AV-HuBERT on the Violin dataset showed that these models performed poorly, yielding a 78.3% WER. This failure is primarily due to the inconsistent face resolution and frequent off-screen speakers inherent in TV series, which prevent these models from consistently finding or resolving the lip-movement cues they rely on.
We didn't include AV-HuBERT as a baseline because it relies on high-resolution, aligned face tracks to function, meaning that it is ill-suited for this complex domain. Consequently, we focus our comparison on post-ASR correction methods that leverage high-level semantic context, which is more robust to these visual challenges.
@@ 5.3 Results Analysis
Figure 2: Comparison between results generated by WavLM and its corresponding variant augmented with VPC.
5.4 Prompt Sensitivity Analysis
A common concern with Large Multimodal Models (LMMs) and LLMs is their sensitivity to prompt engineering, which can lead to ``hallucinations" or inconsistent performance. To evaluate the robustness of our VPC framework, we conducted a pilot study on a randomly selected 100-clip subset of the Violin dataset to evaluate three prompt strategies, we conducted a sensitivity analysis by testing three different prompt templates for the VLMM: (1) Coarse-QA, asking for the name of the show and provide background info, (2) Fine-QA, asking for more detailed context of the shown clip, include scenes, actor (3) All-QA (chosen) that asks both Coarse and Fine QA.
Table 2: Prompt sensitivity analysis (WER %) on a 100-clip pilot subset.
Prompt Strategy Wav2Vec 2.0 HuBERT WavLM
Coarse-QA 25.32 22.90 23.37
Fine-QA 25.41 23.15 23.72
All-QA (Chosen) 25.30 22.93 22.54
As shown in Table 2, we observed marginal changes in WER across different prompt strategies, indicating the framework's robustness to specific phrasing. Interestingly, without the broader scene context, the performance of Fine-QA (detailed descriptions) was slightly inferior to Coarse-QA (overall context). Consequently, we selected the All-QA strategy in the main results for all subsequent experiments, as it effectively captures both high-level situational context and fine-grained visual details, providing the most stable performance gains across backbones.
@@ 5.5 Case Study
We further show a case study about the comparison between results generated by WavLM and its corresponding variant augmented with our method. In this sample, our proposed model revises the error words ``a be hi hat'' into the correct word ``a beehive''. This showcases ability of the context-dependent disambiguation since the video content provides additional information to facilitate the LLM to revise the potential error by reasoning. This further verifies the effectiveness of our method.
@@ 6 Conclusion
7 Generative AI Use Disclosure
During the preparation of this work, all (co-)authors only used Gen AI tools to review and make corrections on grammar and choice of words. After the editing and polishing, the (co-)authors reviewed and edited the content as needed and take full responsibility for the content of the paper.